If you have 12 eggs, 3 of which are cracked, and you choose 5, what is the probability of 1) choosing all 3 cracked 2) choosing no cracked and 3) choosing 2 cracked.
Answer by stanbon(57323)

Binomial Problem with:
n = 5 and p(cracked) = 1/4
---
If you have 12 eggs, 3 of which are cracked, and you choose 5, what is the probability of
1) P(choosing all 3 cracked) = 5C3(1/4)^3(3/4)^2 = 0.08789
------
2) P(choosing no cracked) 5C0(1/4)^0*(3/4)^5 = 0.2373
-------------------------------
3) P(choosing 2 cracked) = 5C2(1/4)^2(3/4)^3 = 0.2637
=======================================================
Cheers,
Stan H.